At year-end 2016, total assets for Arrington Inc. were $1.8 million and accounts payable were $300,000. Sales, which in 2016 were $2.1 million, are expected to increase by 15% in 2017. Total assets and accounts payable are proportional to sales, and that relationship will be maintained; that is, they will grow at the same rate as sales. Arrington typically uses no current liabilities other than accounts payable. Common stock amounted to $445,000 in 2016, and retained earnings were $205,000. Arrington plans to sell new common stock in the amount of $60,000. The firm's profit margin on sales is 6%; 45% of earnings will be retained.

  1. What were Arrington's total liabilities in 2016? Write out your answer completely. For example, 25 million should be entered as 25,000,000. Round your answer to the nearest cent.
    $

  2. How much new long-term debt financing will be needed in 2017? Write out your answer completely. For example, 25 million should be entered as 25,000,000. Do not round your intermediate calculations. Round your answer to the nearest cent. (Hint: AFN - New stock = New long-term debt.)
    $
Answer #1

(a) Total Assets = $ 1.8 million or $ 1800000, Accounts Payable = Current Liabilities = $ 300000 (Accounts Payable are the only Current Liabilities being used), Common Stock = $ 445000, Retained Earnings = $ 205000,

Total Assets = Common Stock + Retained Earnings + Total Liabilities

Total Liabilities = 1800000 - 445000 - 205000 = $ 1150000

Current Liabilities = $ 300000

Long-Term Debt = 1150000 - 300000 = $ 850000

(b) 2016 Sales = $ 2.1 million, Expected Sales Increase = 15 %

2017 Sales = 2100000 x 1.15 = $ 2415000

As total assets and accounts payable are both directly proportional to sales, these two parameters should also increase by 15 %

Expected Total Assets 2017 = 1800000 x 1.15 = $ 2070000

Expected Current Liabilities 2017 = 300000 x 1.15 = $ 345000

Net Profit Margin = 6% of Expected Sales 2017, Net Profit 2017 = 0.06 x 2415000 = $ 144900

Addition to Retained Earnings = 45 % of Net Income = 0.45 x 144900 = $ 65205

Additional Financing Needed (AFN) = Total Assets 2017 - Common Stock - Retained Earnings - Addition to Retained Earnings - Long-Term Debt - Current Liabilities 2017 = 2070000 - 445000 - 205000 - 65205 - 850000 - 345000 = $ 159795

New Stock Issued = $ 60000

AFN - New Stock Issued = New Long-Term Debt

New Long-Term Debt = 159795 - 60000 = $ 99795